# Quantum-Enhanced Neural Exchange-Correlation Functionals

### v2

@@ IV Results and discussion
Figure 4: Results of training neural XC functionals on different geometries of hydrogen molecule in 1D KS-DFT framework Li et al. (2021). The grey area highlights results within chemical precision ($\leq 1.6$ mHa) and vertical dashed lines mark the training data. Local models improve upon the LDA shape, but still produce a poor qualitative agreement with the reference. Best results are obtained for global models when the whole density is considered as model input. The XC models are defined in Section II.3. All neural functionals are trained on eight training data points (circles) and the final model parameters are selected according to the lowest loss on the validation results. The reference (Ref) results are generated using DMRG by Li et al. Li et al. (2021)
In this section, we present the outcomes of numerical experiments and comment on their performance in 1D and 3D settings. First, we evaluate all our models on H2 using the 1D KS-DFT framework based on the implementation of Li et al. Li et al. (2021) Next, we extend our approach to the 3D framework based on PySCFAD Zhang and Chan (2022) and test the extrapolation capacity of our better models on the H2 molecule. We assess the effect of the fractional orbital occupation on H2 and evaluate our best model on planar H4. We evaluate the generalization capability on the H2H2 system, which was not part of the training dataset, in a 1D setting. Finally, we estimate the number of measurements required in a quantum hardware implementation. All calculations are performed on a classical computer using Qadence. Seitz et al. (2024) For all our results, we define the density error in terms of the mean squared error (MSE) and include the local-density approximation (LDA) functional to contrast with our ML-based functionals. The hyperparameter search is performed for all proposed models and the best parameter combination (e.g., number of qubits $N_{q}$, layers $N_{l}$ and depth $N_{d}$) is selected.
@@ IV.1 Local versus global functionals
In order to evaluate if local quantum models can improve on the Local MLP, we propose the Local ChQNN and PrQNN. These models employ the Chebyshev and product feature maps (see Appendix C), respectively. We use $N_{q}=4$ qubits, $N_{l}=1$ (no data reuploading) and $N_{d}=8$ layers for our ansatz. The results in terms of non-parallelity error (NPE) of the energy are very similar to those of the Local MLP model. The energy and density profiles of local quantum and classical models are qualitatively similar, as expected due to locality of both models. It is therefore necessary to verify if including more information in the quantum models brings benefits in terms of lower energy error $|\Delta E|$ and MSE of the density profile.
For that purpose, we evaluate a global version of MLP with $N_{\mathrm{grid}}=513$ input neurons. This allows the model to consider the whole density to infer the XC energy via the global embedding, hence there is only one output neuron. This results in $\sim 2.6\cdot 10^{5}$ parameters as reported in Table 1. The corresponding energy profile is significantly improved (lower error with respect to the reference) in comparison to all other models. Similarly, the proposed KSR model Li et al. (2021) can be tailored to any locality. The global version of KSR model was shown to perform best and generalize perfectly in a single molecule even from only 2 data points (here 8). We used the same density initialization as for Global MLP and the model struggled to generalize correctly, leading to MSE on density comparable to LDA, despite rendering chemically accurate energies. However, KSR is significantly more parameter-efficient, having only $\sim 10^{3}$ parameters versus $\sim 2.6\cdot 10^{5}$ for Global MLP, which are generally known to overfit the training data. In addition, given good initialization, global KSR is capable of learning the dissociation profile of H2 perfectly only from 2 data points, as shown in ref. Li et al., 2021. For all the remaining models with global embedding, the results perform as well or even better than local LDA-like functionals, at least for the total energies. Moreover, the global models have the lowest non-parallelity error (NPE) by an order of magnitude in comparison to local models (see Tab. 1). A significant reduction (3-5 orders of magnitude) in average MSE of the electron density is achieved only by fully global models such as Global QNN and Global MLP.
@@ IV.3 Hybrid and quantum-inspired quantum-classical models
Since local QNNs are not capable of correctly fitting even the training data, we explore architectures that allow us to embed $L$ data points and hence approach non-local functionals. In particular, in the Global QNN model, we consider the idea of embedding $L$ data points in a single QNN with $N_{q}=6$ qubits, $N_{l}=1$ (no data reuploading) and $N_{d}=8$ layers, and apply the same model onto $N_{\mathrm{grid}}/L$ different batches. The results are combined via a single-layer MLP with 171 input neurons and 1 output for the global embedding. This allows us to embed chunks of data into a limited quantum model with a restricted number of qubits. By combining such outputs, we obtain a fully global functional. The model has only $\sim 100$ trainable parameters, which is orders of magnitude smaller than for the Global MLP, since both are composed of most general quantum/classical layers. Note that comparing it with the KSR model is not reasonable because the latter possesses numerous inductive biases such as the self-interaction gate and the softly imposed negativity of XC energy.
We also evaluated the amplitude embedding as a feature map, using a DQC with $N_{q}=9$ qubits, $N_{l}=1$, $N_{d}=10$. We name this model Global QiNN since, in general, it is hard to embed the classical data (electron density) exactly in a polynomial number of one-/two-qubit gates. Park et al. (2019) The number of qubits was chosen to fit all the elements of $\bm{n}$ exactly in a state-vector. In Tab. 1, the results for Global QiNN are in good agreement with the reference: within chemical precision and overall density error that significantly improved upon LDA. Notably, the outcome is similar to Global QNN, which is simple to implement on near-term devices. The Global QiNN architecture, despite producing satisfactory results, does not produce a significant advantage over quantum-classical hybrid QNN. It can, however, serve for theoretical studies since only quantum layers are employed, which dissociates it completely from classical models.
@@ IV.4 Extension from 1D to 3D KS-DFT framework
To demonstrate the ability to efficiently train quantum XC models in our 3D framework, we evaluate our most complex neural functional, Global QiCNN, in such a setting. Similarly to the 1D version, the model consists of 3 chained quantum layers, each with $N_{q}=4$ qubits, $N_{l}=1$ and $N_{d}=14$ ansatz layers. The final quantum layer connects to MLP with 149 neurons to output the total XC energy. As depicted in Fig. 5, Global QiCNN suffers from energy oscillations and is unable to extrapolate correctly. This is the ideal situation to demonstrate the ability of fractional occupation (FO) methods to alleviate such problems, which are apparent in 3D due to increased complexity. To that end, we show that even with 3 data points (of which 2 are almost at the same bond distance) the complex model can be regularized in order to fit better the dissociation curve, e.g., lower NPE (see Tab. 2). Despite not being within chemical precision along the whole energy profile, the density error is improved by orders of magnitude in comparison to the LDA result. This is achieved with a FO algorithm Chai (2012) that is discussed in Appendix A. The FO approach relaxes the constraint of the integer orbital occupation in PySCF’s restricted KS-DFT implementation and allows us to approximate the true fractional ones. In addition, the neural XC functional is optimized through training, thereby compensating for errors arising from the approximation of orbital occupations. As reported in Tab. 2, the use of the FO approach results in an approximately 5-fold reduction in both the average energy error and its variance, without any extra model parameters. Note that the FO method is not limited to use only with quantum models; it can also be applied more broadly for training classical neural XC functionals.
@@
As the ultimate test of our 3D framework, demonstrating the capability to accurately learn from challenging molecular systems, we trained our best hybrid architecture, the Global QNN, on planar H4. This is a typical small system exhibiting strong correlation, commonly used to benchmark ab initio methods. Bulik et al. (2015); Sokolov et al. (2020) Here, Global QNN consists of one quantum layer with $N_{q}=4$ qubits, $N_{l}=1$, $N_{d}=14$ and followed by MLP with 596 neurons that outputs $E_{\mathrm{XC}}$. In Fig. 6, we note that even after the shift, the LDA approximation greatly overestimates the energy barrier at $\beta=90^{\mathrm{o}}$. This quantum-classical hybrid model, given its limited expressivity, provides the best achievable fit of the training data and extrapolates smoothly to various geometries. In Tab. 3, we note that the model achieves chemical precision over the whole energy profile (NPE $\approx 10^{-4}$). Concerning density errors, they are of the order of $10^{-5}$, similar to the H2 density errors (Tab. 2) that are obtained within the same 3D framework. Hence, we demonstrated that provided accurate reference data, it is possible to train quantum-enhanced functionals to learn from strongly correlated systems. In the future, it would be interesting to evaluate if such models can generalize correctly from small to other larger strongly correlated systems such as stretched N2.
@@
Figure 6: Extrapolation of Global QNN (Eq. (18)) on planar H4 within the 3D framework with a fractional occupation approach. Chai (2012) Global QNN achieves chemical precision in energy, surpassing the LDA results by orders of magnitude. In the top panel, LDA and Global QNN results are shifted to match the reference data at extremities to compare the shapes of the dissociation profiles. Shifts can be inferred from the $|\Delta E|$ panel. The reference (Ref) results are generated with the FCI/6-31G level of theory.
@@ IV.5 Zero-shot generalization of XC functionals
Figure 7: Generalization of Global QNN (Eq. (18)) to unseen H2H2 system trained on H2 and H4 molecules within the 1D framework. The model achieves chemical precision over some regions of the dissociation profile without being explicitly trained on it. The reference (Ref) results are generated using DMRG by Li et al. Li et al. (2021)
Our results highlight Global QNN as the simplest hybrid model, comprising a single quantum and classical layer, capable of achieving chemical precision for H2. For its training, the dataset only contained the hydrogen molecule at different geometries. Such a dataset does not allow the model to implicitly learn the XC energy dependency on the variable number of electrons $N_{e}$ in the input density (i.e., $N_{e}=\sum_{i=1}^{N_{\mathrm{grid}}}w_{i}{n}_{i}$) when evaluating on a different system such as H2H2 (separate hydrogen molecules). Hence, we enrich our train dataset with molecules such as linear H4 at different bond distances, as done in ref. Li et al. (2021) We train Global QNN using the following bond distances: For H2, $R\in(0.48,1.28,3.04,3.84)$ Bohr and for H4, $R\in(1.28,2.08,3.36,4.48)$ Bohr.
In Fig. 7, we note that chemical precision is achieved only for small values of $R<2$ Bohr and momentarily at $R=5.3$ Bohr, with the overall energy closely aligned with the reference, resulting in lower energy error compared to LDA. MSE on the density consistently remains lower than the LDA result. The XC model yields suboptimal qualitative predictions, evidenced by the energy minimum at approximately $3.5$ Bohr (stable configuration). Considering the utilization of only basic classical and quantum layers and a few data points, these results demonstrate the capacity of our models to generalize. In fact, as stated in Tab. 4, our model has only 144 trainable parameters compared to the state-of-the-art Global KSR model, which has 10 times more. Hence, addressing the shortcomings would require improving the architecture of our model (e.g., using molecular symmetries) and expanding the training dataset, which is beyond the scope of this study.
Table 4: Zero-shot generalization results for Global QNN evaluated on H2H2, trained on H2 and H4 within the 1D KS-DFT framework (Fig. 7).
@@
$N_{q}$/$N_{d}$/$N_{l}$ 6/8/1
No. parameters 144
H2H2 Av. $|\Delta E|$ [Ha] $(4.9\pm 1.4)\cdot 10^{-2}$
@@
Figure 8: Dependence of the average energy error on the size of classical and quantum (hybrid) models. Quantum models are consistently more compact ($\leq 10^{3}$ parameters) than classical models (e.g., GKSR with $1568$ parameters) given a fixed energy error, while being still capable of achieving chemical precision.

### v3

@@ IV Results and discussion
Figure 4: Results of training neural XC functionals on different geometries of hydrogen molecule in 1D KS-DFT framework ksr. The grey area highlights results within chemical precision ($\leq 1.6$ mHa) and vertical dashed lines mark the training data. Local models improve upon the LDA shape, but still produce a poor qualitative agreement with the reference. Best results are obtained for global models when the whole density is considered as model input. The XC models are defined in Section II.3. All neural functionals are trained on eight training data points (circles) and the final model parameters are selected according to the lowest loss on the validation results. The reference (Ref) results are generated using DMRG by Li et al. ksr
In this section, we present the outcomes of numerical experiments and comment on their performance in 1D and 3D settings. First, we evaluate all our models on H2 using the 1D KS-DFT framework based on the implementation of Li et al. ksr Next, we extend our approach to the 3D framework based on PySCFAD zhang2022differentiable and test the extrapolation capacity of our better models on the H2 molecule. We assess the effect of the fractional orbital occupation on H2 and evaluate our best model on planar H4. We evaluate the generalization capability on the H2H2 system, which was not part of the training dataset, in a 1D setting. Finally, we estimate the number of measurements required in a quantum hardware implementation. All calculations are performed on a classical computer using Qadence. qadence2024pasqal
For all our results, we define the density error in terms of the mean squared error (MSE) and include the local-density approximation (LDA) functional to contrast with our ML-based functionals. In addition, we analyze the results in terms of the non-parallelity error (NPE) tuna2015assessment, defined as the absolute difference between the largest and smallest energy errors $\Delta E$ over the whole energy profile. The average $|\Delta E|$ measures the overall closeness to the true dissociation profile, whereas the NPE measures non-parallelism after removing any constant shift. In reaction rate calculations, the absolute energy scale is irrelevant; only the relative shape of the energy profile matters, e.g., see ref. douroudgari2021atmospheric. For example, one could have $\text{average}~|\Delta E|=1$ Ha (constant shift) but $\text{NPE}<1$ mHa, indicating that the correct chemical behavior (e.g., reaction rates) is preserved, enabling accurate predictions. Regarding the MSE on the density, in ref. kim2013understanding it was shown that, in typical cases, density errors do not significantly affect the total energy, except in specific scenarios. Therefore, in practical applications, NPE is often the most important metric.
The hyperparameter search is performed for all proposed models and the best parameter combination (e.g., number of qubits $N_{q}$, layers $N_{l}$ and depth $N_{d}$) is selected.
@@ IV.1 Local versus global functionals
In order to evaluate if local quantum models can improve on the Local MLP, we propose the Local ChQNN and PrQNN. These models employ the Chebyshev and product feature maps (see Appendix C), respectively. We use $N_{q}=4$ qubits, $N_{l}=1$ (no data reuploading) and $N_{d}=8$ layers for our ansatz. The results in terms of NPE of the energy are very similar to those of the Local MLP model. The energy and density profiles of local quantum and classical models are qualitatively similar, as expected due to locality of both models. It is therefore necessary to verify if including more information in the quantum models brings benefits in terms of lower energy error $|\Delta E|$ and MSE of the density profile.
For that purpose, we evaluate a global version of MLP with $N_{\mathrm{grid}}=513$ input neurons. This allows the model to consider the whole density to infer the XC energy via the global embedding, hence there is only one output neuron. This results in $\sim 2.6\cdot 10^{5}$ parameters as reported in Table 1. The corresponding energy profile is significantly improved (lower error with respect to the reference) in comparison to all other models. Similarly, the proposed KSR model ksr can be tailored to any locality. The global version of KSR model was shown to perform best and generalize perfectly in a single molecule even from only 2 data points (here 8). We used the same density initialization as for Global MLP and the model struggled to generalize correctly, leading to MSE on density comparable to LDA, despite rendering chemically accurate energies. However, KSR is significantly more parameter-efficient, having only $\sim 10^{3}$ parameters versus $\sim 2.6\cdot 10^{5}$ for Global MLP, which are generally known to overfit the training data. In addition, given good initialization, global KSR is capable of learning the dissociation profile of H2 perfectly only from 2 data points, as shown in ref. ksr. For all the remaining models with global embedding, the results perform as well or even better than local LDA-like functionals, at least for the total energies. Moreover, the global models have the lowest non-parallelity error (NPE) by an order of magnitude in comparison to local models (see Tab. 1). A significant reduction (3-5 orders of magnitude) in average MSE of the electron density is achieved only by fully global models such as Global QNN and Global MLP.
@@ IV.3 Hybrid and quantum-inspired quantum-classical models
Since local QNNs are not capable of correctly fitting even the training data, we explore architectures that allow us to embed $L$ data points and hence approach non-local functionals. In particular, in the Global QNN model, we consider the idea of embedding $L=3$ data points in a single QNN with $N_{q}=6$ qubits, $N_{l}=1$ (no data reuploading) and $N_{d}=8$ layers, and apply the same model onto $N_{\mathrm{grid}}/L$ different batches. The results are combined via a single-layer MLP with 171 input neurons, one bias parameter, and one output for the global embedding. This allows us to embed chunks of data into a limited quantum model with a restricted number of qubits. By combining such outputs, we obtain a fully global functional. The model has only $\sim 300$ trainable parameters, which is orders of magnitude smaller than for the Global MLP, since both are composed of most general quantum/classical layers. Note that comparing it with the KSR model is not reasonable because the latter possesses numerous inductive biases such as the self-interaction gate and the softly imposed negativity of XC energy.
We also evaluated the amplitude embedding as a feature map, using a DQC with $N_{q}=9$ qubits, $N_{l}=1$, $N_{d}=10$. We name this model Global QiNN since, in general, it is hard to embed the classical data (electron density) exactly in a polynomial number of one-/two-qubit gates. park2019circuit The number of qubits was chosen to fit all the elements of $\mathbf{n}$ exactly in a state-vector. In Tab. 1, the results for Global QiNN are in good agreement with the reference: within chemical precision and overall density error that significantly improved upon LDA. Notably, the outcome is similar to Global QNN, which is simple to implement on near-term devices. The Global QiNN architecture, despite producing satisfactory results, does not produce a significant advantage over quantum-classical hybrid QNN. It can, however, serve for theoretical studies since only quantum layers are employed, which dissociates it completely from classical models.
@@ IV.4 Extension from 1D to 3D KS-DFT framework
To demonstrate the ability to efficiently train quantum XC models in our 3D framework, we evaluate our most complex neural functional, Global QiCNN, in such a setting. Similarly to the 1D version, the model consists of 3 chained quantum layers, each with $N_{q}=4$ qubits, $N_{l}=1$ and $N_{d}=14$ ansatz layers. The final quantum layer connects to MLP with 149 neurons to output the total XC energy. As depicted in Fig. 5, Global QiCNN suffers from energy oscillations and is unable to extrapolate correctly. This is the ideal situation to demonstrate the ability of fractional occupation (FO) methods to alleviate such problems, which are apparent in 3D due to increased complexity. To that end, we show that even with 3 data points (of which 2 are almost at the same bond distance) the complex model can be regularized in order to fit better the dissociation curve, e.g., lower NPE (see Tab. 2). Despite not being within chemical precision along the whole energy profile, the density error is improved by orders of magnitude in comparison to the LDA result. This is achieved with a FO algorithm frac-occ that is discussed in Appendix A. The FO approach relaxes the constraint of the integer orbital occupation in PySCF’s restricted KS-DFT implementation and allows us to approximate the true fractional ones. In addition, the neural XC functional is optimized through training, thereby compensating for errors arising from the approximation of orbital occupations. As reported in Tab. 2, the use of the FO approach results in an approximately 5-fold reduction in both the average energy error and its variance, without any extra model parameters. Note that the FO method is not limited to use only with quantum models; it can also be applied more broadly for training classical neural XC functionals.
@@
As the ultimate test of our 3D framework, demonstrating the capability to accurately learn from challenging molecular systems, we trained our best hybrid architecture, the Global QNN, on planar H4. This is a typical small system exhibiting strong correlation, commonly used to benchmark ab initio methods. bulik2015can; sokolov2020quantum Here, Global QNN consists of one quantum layer with $N_{q}=4$ qubits, $N_{l}=1$, $N_{d}=14$ and followed by MLP with 596 neurons that outputs $E_{\mathrm{XC}}$. In Fig. 6, we note that the LDA approximation greatly overestimates the energy barrier at $\beta=90^{\mathrm{o}}$. This quantum-classical hybrid model, given its limited expressivity, provides the best achievable fit of the training data and extrapolates smoothly to various geometries. In Tab. 3, we note that the model achieves chemical precision over the whole energy profile (NPE $\approx 10^{-4}$). Concerning density errors, they are of the order of $10^{-5}$, similar to the H2 density errors (Tab. 2) that are obtained within the same 3D framework. Hence, we demonstrated that provided accurate reference data, it is possible to train quantum-enhanced functionals to learn from strongly correlated systems. In the future, it would be interesting to evaluate if such models can generalize correctly from small to other larger strongly correlated systems such as stretched N2.
@@
Figure 6: Extrapolation of Global QNN (Eq. (18)) on planar H4 within the 3D framework with a fractional occupation approach. frac-occ Global QNN achieves chemical precision in energy, surpassing the LDA results by orders of magnitude. In the top panel, LDA and Global QNN results are shifted to match the reference data at extremities to compare the shapes of the dissociation profiles. Shifts can be inferred from the $|\Delta E|$ panel. The reference (Ref) results are generated with the FCI/6-31G level of theory.
@@ IV.5 Zero-shot generalization of XC functionals
Figure 7: Generalization of Global QNN (Eq. (18)) to unseen H2H2 system trained on H2 and H4 molecules within the 1D framework. The model achieves chemical precision over some regions of the dissociation profile without being explicitly trained on it. The reference (Ref) results are generated using DMRG by Li et al. ksr
Our results highlight Global QNN as the simplest hybrid model, comprising a single quantum and classical layer, capable of achieving chemical precision for H2. For its training, the dataset only contained the hydrogen molecule at different geometries. Such a dataset does not allow the model to implicitly learn the XC energy dependency on the variable number of electrons $N_{e}$ in the input density (i.e., $N_{e}=\sum_{i=1}^{N_{\mathrm{grid}}}w_{i}{n}_{i}$) when evaluating on a different system such as H2H2 (separate hydrogen molecules). Hence, we enrich our train dataset with molecules such as linear H4 at different bond distances, as done in ref. ksr We train Global QNN using the following bond distances: For H2, $R\in(0.48,1.28,3.04,3.84)$ Bohr and for H4, $R\in(1.28,2.08,3.36,4.48)$ Bohr.
Table 4: Zero-shot generalization results for Global QNN evaluated on $H2$, trained on H2 and H4 within the 1D KS-DFT framework (Fig. 7).
@@
$N_{q}$$N_{d}$$N_{l}$ 6/8/1
No. parameters 316
$H2$ Av. $|\Delta E|$ [Ha] $(4.9\pm 1.4)\cdot 10^{-2}$
@@
In Fig. 7, we note that chemical precision is achieved only for small values of $R<2$ Bohr and momentarily at $R=5.3$ Bohr, with the overall energy closely aligned with the reference, resulting in lower energy error compared to LDA. MSE on the density consistently remains lower than the LDA result. The XC model yields suboptimal qualitative predictions, evidenced by the energy minimum at approximately $3.5$ Bohr (stable configuration). Considering the utilization of only basic classical and quantum layers and a few data points, these results demonstrate the capacity of our models to generalize. In fact, as stated in Tab. 4, our model has only 316 trainable parameters compared to the state-of-the-art Global KSR model, which has $\approx 6$ times more. Hence, addressing the shortcomings would require improving the architecture of our model (e.g., using molecular symmetries) and expanding the training dataset, which is beyond the scope of this study.